OpenStudy (gabylovesyou):

(-1, 3) and perpendicular to y = -1/3x + 7. y = 3x - 12 y = 3x + 6 y = 3x - 6 y = 3x

OpenStudy (anonymous):

i might as well just teach you everything about lines......

OpenStudy (anonymous):

parallel lines share the same slope perpendicular lines slopes are negative reciprocals of each other ex. if one line had a slope of 2, the perpendicular line would have a slope of -1/2 Point slope form \[y-y_1=m(x-x_1)\] when ever you are trying to find the equation of a line and you're given the slope and a point, substitute those into the point slope form then solve for y to get slope intercept form \[y=mx+b\] where m is the slope and b is the y intercept

OpenStudy (gabylovesyou):

what do i do to start??

OpenStudy (anonymous):

find the slope of the perpendicular line

OpenStudy (gabylovesyou):

the slope ofy = -1/3x + 7 ???? its -1/3

OpenStudy (anonymous):

so the negative reciprocal of -1/3 is?

OpenStudy (gabylovesyou):

3?

OpenStudy (anonymous):

so the slope of your perpendicular line is 3 now you have a point and a slope, substitute those into the point slope form

OpenStudy (gabylovesyou):

what do i substitute with what???????????

OpenStudy (anonymous):

(-1, 3) and m=3 into y-y1...

OpenStudy (gabylovesyou):

mmmmmm

OpenStudy (anonymous):

he's referring to the point slope form. \[y-y1=m(x-x1)\] We now know that m=3, x1=-1 and y1=3 \[y-3=3(x+1)\] Solve for y.

OpenStudy (gabylovesyou):

y = 3?

OpenStudy (anonymous):

\[y=3x+3+3\]

OpenStudy (anonymous):

what do you get?

OpenStudy (gabylovesyou):

y = 3 (x + 2)

OpenStudy (anonymous):

Yes. But what is your answer?

OpenStudy (gabylovesyou):

y = 3x + 6

OpenStudy (anonymous):

yes.

OpenStudy (gabylovesyou):

thank you

OpenStudy (anonymous):

you're welcome.